50 grams of CaCo3 is allowed to react with 73.5 grams of h3po4 calculate amount of Ca3(Po4)2 formed?

Grade:11

4 Answers

DEEPTHI JANGA
7 years ago
good evening yasik guptha,
in my point of veiw,
balanced equation- 3CaCo3+2H2Po4 gives Ca3(Po4)2+3Co2+3H2O  so
210g of CaCo3 and 196 H3Po4 gives 310g of Ca3(Po4)
According to problem,
50gof CaCo3 and 73.5 gH3Po4 gives x g of  Ca3(pO4) then ‘
x=(50+73.5)*310/210+196
i.e. equal to 75.66205g
Kshitij nandwana
6 years ago
I th
anonymous
5 years ago
applying poac...
let ca be conserved..
so we have
1×moles of caco3=3×moles of ca3(po4)3
1/2=3×moles of ca3(po4)3
moles=1/6
now wt=1/6×310
Aashay
5 years ago
3 mole of CaCO3 -: 1 mole Ca3(PO4)2
1 mole of CaCO3 -: 1/3 mole Ca3(PO4)2
1/2 mole CaCO3 -: 1/6 mole Ca3(PO4)2
So 1/6 moles of Ca3(PO4)2 will have mass 51.666 g